OpenStudy (geekfromthefutur):

Help plz i will give medal and fan my sister cant figure this out and I need some help solvin Part C: Another object moves in the air along the path of g(t) = 28 + 48.8t where g(t) is the height, in feet, of the object from the ground at time t seconds. Use a table to find the approximate solution to the equation H(t) = g(t), and explain what the solution represents in the context of the problem? [Use the function H(t) obtained in Part A, and estimate using integer values] (4 points) Part D: Do H(t) and g(t) intersect when the projectile is going up or down, and how do you know?

OpenStudy (geekfromthefutur):

pls help me

OpenStudy (geekfromthefutur):

@Kainui @Loser66

OpenStudy (loser66):

You must post the whole problem since part C links to part A which says what H(t) is

OpenStudy (geekfromthefutur):

I said for part a Part A: H(t) = -16t^ 2 + vt + s H(t) = -16t^ 2 + (50)t + (90) H(t) = -16t^ 2 + 50t + 90

OpenStudy (geekfromthefutur):

The function H(t) = -16t2 + vt + s shows the height H (t), in feet, of a projectile launched vertically from s feet above the ground after t seconds. The initial speed of the projectile is v feet per second. Part A: The projectile was launched from a height of 90 feet with an initial velocity of 50 feet per second. Create an equation to find the time taken by the projectile to fall on the ground. (2 points)

OpenStudy (geekfromthefutur):

@Loser66

OpenStudy (loser66):

if they say "use the table..." then, you make a table.

OpenStudy (loser66):

that is 

OpenStudy (loser66):

to see when they are same

OpenStudy (geekfromthefutur):

Oh ok

OpenStudy (loser66):

for part c) just let H(t) =g(t) and solve for t

OpenStudy (geekfromthefutur):

Dont mean to put extra work on ya but how do you let them solve for t

OpenStudy (loser66):

the graphs intersect mean they are equal at some points, maximum 2 points. H(t) is a downward parabola g(t) is a line with positive slope, they have graphs like

OpenStudy (loser66):